For N2(g) + 3 H2(g) ↔ 2 NH3(g), Kp = 4.34 × 10-3 at 300℃. What is Kp or the reverse reaction?
1
Expert's answer
2020-10-23T09:16:34-0400

Kp for reverse reaction = 1/Kp = 1/4.34*10^-3 = 230,41